# Crowd-PrefRL: Preference-Based Reward Learning from Crowds

### v1

@@ Abstract
Preference-based reinforcement learning (RL) provides a framework to train agents using human feedback through pairwise preferences over pairs of behaviors, enabling agents to learn desired behaviors when it is difficult to specify a numerical reward function. While this paradigm leverages human feedback, it currently treats the feedback as given by a single human user. Meanwhile, incorporating preference feedback from crowds (i.e. ensembles of users) in a robust manner remains a challenge, and the problem of training RL agents using feedback from multiple human users remains understudied. In this work, we introduce Crowd-PrefRL, a framework for performing preference-based RL leveraging feedback from crowds. This work demonstrates the viability of learning reward functions from preference feedback provided by crowds of unknown expertise and reliability. Crowd-PrefRL not only robustly aggregates the crowd preference feedback, but also estimates the reliability of each user within the crowd using only the (noisy) crowdsourced preference comparisons. Most importantly, we show that agents trained with Crowd-PrefRL outperform agents trained with majority-vote preferences or preferences from any individual user in most cases, especially when the spread of user error rates among the crowd is large. Results further suggest that our method can identify minority viewpoints within the crowd.
@@ 1 Introduction
Figure 1: Crowd-PrefRL framework for preference-based RL from crowds. We assume a crowd, or ensemble, of users is queried for their preferences over pairs of segmented behaviors (A vs. B). These crowd preferences are then used to learn a reward function for preference-based RL.
Reinforcement learning (RL) from human feedback Casper et al. (2023) is a promising approach for learning intelligent behaviors in the absence of a known numerical reward signal. Such methods have shown success in domains such as Atari games Christiano et al. (2017), robotics Wilde et al. (2021); Torne et al. (2023); Lee et al. (2021), and large language models (LLMs) Ouyang et al. (2022). In this work, we consider human feedback that is crowdsourced, i.e., where (1) data is labeled by multiple human users (e.g. $\geq 2$ labels for each data point), and (2) these labels are aggregated to form an ensemble label to be used for downstream analysis. In applications from LLMs to household robotics to autonomous driving, crowdsourced human data could enable RL algorithms to learn user-preferred behaviors that are robust across many real-world situations. Yet, at the same time, crowdsourcing algorithms will need to recognize and account for disagreement across users in the crowd. Example scenarios could include: an algorithm that a) trains models that are sensitive to multiple user viewpoints (e.g., an LLM provides information that benefits an underrepresented minority at minimal cost to the majority Siththaranjan, Laidlaw, and Hadfield-Menell (2023)), b) trains models that reflect a majority viewpoint while enabling downstream individual personalization (e.g., a household robot demonstrates performant out-of-the-box behavior, but fine-tunes to specific users’ preferences for tidying and organization), and c) identifies and rejects malicious user data (e.g., an LLM is resilient to feedback containing misinformation).
This work considers RL from crowdsourced pairwise preferences, in which a human answers queries of the form “Do you prefer A or B?” Christiano et al. (2017); Lee, Smith, and Abbeel (2021); Lee et al. (2021). Current preference-based RL methods treat preference feedback as given by a single human user, even if the feedback is actually given by multiple users with potentially different backgrounds and levels of expertise. For example, when learning reward functions for training LLMs, each individual preference pair query is typically shown to just one of the users due to multiple constraints (e.g. cost, human labeling time/effort); thus, the full labeled dataset aggregates the disjoint query subsets shown to each user Ouyang et al. (2022). However, it may often be preferable to not only obtain feedback from a group of users, but also to aggregate the crowdsourced data in a non-naive manner that recognizes differences between individuals. This capability can be beneficial for multiple reasons, for example to analyze patterns of disagreement in preference feedback across the crowd and to detect minority groups in the crowd (i.e., users that provide feedback that is consistently different from the majority). It also holds important practical modeling implications, for example mitigating model bias and ensuring fairness Ouyang et al. (2022) and assisting users with diverse preferences or varying degrees of noise in their feedback Siththaranjan, Laidlaw, and Hadfield-Menell (2023).
While previous literature uses crowdsourcing methods to aggregate labels across users and learn crowd preferences Chen et al. (2013); Li et al. (2016), the specific application of crowdsourcing to reward learning in RL remains understudied. Early works that explicitly model human feedback from crowds have focused on imitation learning, where there are multiple demonstrations (assumed to be from multiple human users) of unknown expertise and quality, and the goal is to learn a policy reflecting expert behavior Babes et al. (2011); Dimitrakakis and Rothkopf (2012); Hausman et al. (2017). As noted in Beliaev et al. (2022), however, these approaches simply aggregate the feedback from multiple humans into one homogeneous group, potentially leading to sub-optimal learning due to the presence of varying levels of expertise and reliability among the different human users. Recent work that has explicitly modeled crowdsourced data for imitation learning Beliaev et al. (2022) has shown significant performance gains from simultaneously modeling human demonstrations and estimating the demonstrators’ expertise levels. While recent work by  Zhang and Kashima (2023) used crowdsourcing methods to learn a reward function using preference-based reward learning, it is limited to the offline RL case Levine et al. (2020), where there is a fixed dataset of preference pairs, and the RL algorithm does not have online access to the environment. This potentially limits the expressiveness of the learned reward function to the behaviors contained in the offline dataset.
In this work, we study preference-based RL leveraging feedback from crowds in the online RL scenario, where the RL algorithm has access to the environment and can sample new preference pairs for the crowd to label throughout model training. Specifically:
@@
    We introduce the Crowd-PrefRL framework for performing online preference-based RL leveraging feedback from crowds and instantiate Crowd-PrefRL using Proximal Policy Optimization (PPO) Schulman et al. (2017), a state-of-the-art RL algorithm, to propose Crowd-PrefPPO. Crowd-PrefRL aggregates the user feedback to learn crowdsourced preference labels, and then uses those labels for reward function learning in preference-based RL (see Figure 1).
@@
    We use techniques derived from unsupervised ensemble learning to robustly aggregate the crowd preference feedback, estimating each user’s reliability in a completely unsupervised fashion and detecting the presence of multiple, e.g. minority, viewpoints among the crowd.
  3. 3.
    We perform experiments suggesting that RL agents trained with crowd preferences outperform RL agents trained via majority-vote preferences or preferences from any individual crowd member, in particular when the spread of user errors among the crowd is large. We also show a proof-of-concept for the ability of our method to identify minority viewpoints within a crowd.
This work demonstrates the viability of learning reward functions from preference feedback provided by crowds of unknown expertise and reliability. We believe that this is the first work that focuses on using preference feedback from crowds to learn a crowdsourced reward function for preference-based RL in the online RL setting.
@@ 2 Problem Formulation
In this paper, we are concerned with answering the following questions: (1) How can we effectively combine preference feedback from multiple users of unknown expertise and unknown reliability in the absence of ground-truth information? (2) Can we derive who are the “best” users in a given crowd in an unsupervised manner? (3) Can we learn a crowdsourced reward function that is better than any individual user’s reward function?, and (4) Can we detect the presence of minority feedback (as opposed to purely noisy feedback) in the crowd based only on the preference feedback? This section formulates this problem more precisely.
@@
We consider an agent that takes actions and interacts with an environment Sutton and Barto (2018), such that at each timestep $t$, the agent receives a state $\mathbf{s}_{t}$ from the environment and chooses an action $\mathbf{a}_{t}$ based on its policy $\pi$. In traditional episodic RL, the environment also returns a reward $r(\mathbf{s}_{t},\mathbf{a}_{t})$, and the agent seeks to maximize the total discounted sum of rewards over an episode. However, for many complex tasks, it is difficult to construct a suitable reward function. Preference-based RL approaches instead query a human user to obtain preference feedback over segments of agent behaviors, and use this feedback to learn a reward function  Christiano et al. (2017); Lee, Smith, and Abbeel (2021). A segment is defined as a length-$H$ sequence of observations and actions, $\{(\mathbf{s}_{1},\mathbf{a}_{1}),...,(\mathbf{s}_{H},\mathbf{a}_{H})\}$. Given a pair of segments $x=(A,B)$, a user indicates which segment is preferred: $+1$ in the event that $A$ is preferred over $B$ (denoted by $A\succ B$) or $-1$ if $B$ is preferred over $A$ (denoted by $A\prec B$). Preference-based RL aims to train an agent to perform human-desirable behaviors according to the human user’s preferences.
@@
Our problem formulation assumes access to a crowd, or ensemble, of users, who each assign pairwise preference labels to a sequentially-selected set $\mathcal{S}=\{x_{k}\}_{k=1}^{S}$, $S=|\mathcal{S}|$ of segment pairs. Each query $x_{k}$ takes the form $x_{k}=(A_{k},B_{k})$, where $A_{k},B_{k}$ belong to the set $\mathcal{X}$ of available segments. In addition, we assume that a majority (e.g. at least $51\%$) of the crowd provides preference feedback according to a shared understanding of desired agent behavior. More concretely, we assume the existence of an unknown ground-truth reward $r$, and that for a majority of users, their preferences are (noisily) generated with respect to $r$. Note that this formulation allows a minority group of users to potentially provide feedback according to a different reward function(s). Beyond this, we do not assume any further knowledge about the crowd (e.g. expertise, reliability). Our goal is to learn a crowd-informed preference label for each segment pair using only the preference labels given by the crowd. Formally, let ${\{f_{i}\}_{i=1}^{M}}$ represent $M$ users of unknown reliability and error rate, where each user $i$ provides a preference label $f_{i}(x_{k})\in\{-1,1\}$ on each input query $x_{k}\in\mathcal{S}$. Then, let $\bm{y}$ be the vector of (unknown to the algorithm) ground-truth labels according to the true reward $r$ (i.e., in which segments with higher ground-truth reward are preferred), $\bm{y}=[y_{1},...,y_{S}]^{T}$. Using only the preference labels from $M$ users on the segment pairs in $\mathcal{S}$ and without access to any ground-truth rewards or preferences, we seek an optimal estimate $\hat{\bm{y}}=[\hat{y}_{1},...,\hat{y}_{S}]^{T}$ of the true labels $\bm{y}$ given the crowd preference labels. We also investigate the ability to identify minority groups of users who provide feedback that consistently differs from that of the majority.
@@ 3 Methods
We next explain our methodology for simulating diverse crowds of human users with different rationality levels. Afterwards, we outline our Crowd-PrefRL approach to performing RL with crowdsourced feedback (Algorithm 1): first, we explain how to estimate preference labels from a crowd and then, we show how to integrate these crowd-aggregated preference labels into a framework for preference-based RL.
@@ 3.1 Simulating Diverse Crowds
In order to systematically evaluate and study the effects of user error rates and their impact on learning from crowdsourced user preferences, we apply the Stochastic Preference Model Lee et al. (2021) to simulate users with mixtures of different irrationality levels:
$P[A\succ B;\beta,\gamma]=$ (1)
$\hskip 14.22636pt\frac{e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{i},\mathbf{a}_{t}^{i})\right)}}{e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{i},\mathbf{a}_{t}^{i})\right)}+e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{j},\mathbf{a}_{t}^{j})\right)}},$
where $\gamma\in(0,1]$ is a discount factor to model myopic behavior and $\beta$ is a rationality constant. Note that user feedback is fully rational and deterministic as $\beta\rightarrow\infty$, whereas $\beta=0$ produces uniformly random choices. In addition, we model an additional parameter $\epsilon$ that models users making mistakes due to accidental errors. For our analysis, each user is simulated with a random draw of parameters from the following ranges: $\gamma\in[0.96,1.0]$, $\beta\in[-0.5,-1.0]$, and $\epsilon\in[0.1,0.5]$. These parameter ranges where chosen as they produced diverse user error rates primarily between $15\%$ and $40\%$, which is approximately the range of human error rates observed in a prior study of human feedback in preference-based RL ($12\%-37\%$ in Brown et al. (2019)). We then sample crowds of $M\in\{7,11,15\}$ users according to this model to study the effect of crowd size on crowd label learning.

### v2

@@ Abstract
Preference-based reinforcement learning (RL) provides a framework to train AI agents using human feedback through preferences over pairs of behaviors, enabling agents to learn desired behaviors when it is difficult to specify a numerical reward function. While this paradigm leverages human feedback, it typically treats the feedback as given by a single human user. However, different users may desire multiple AI behaviors and modes of interaction. Meanwhile, incorporating preference feedback from crowds (i.e. ensembles of users) in a robust manner remains a challenge, and the problem of training RL agents using feedback from multiple human users remains understudied. In this work, we introduce a conceptual framework, Crowd-PrefRL, that integrates preference-based RL approaches with techniques from unsupervised crowdsourcing to enable training of autonomous system behaviors from crowdsourced feedback. We show preliminary results suggesting that Crowd-PrefRL can learn reward functions and agent policies from preference feedback provided by crowds of unknown expertise and reliability. We also show that in most cases, agents trained with Crowd-PrefRL outperform agents trained with majority-vote preferences or preferences from any individual user, especially when the spread of user error rates among the crowd is large. Results further suggest that our method can identify the presence of minority viewpoints within the crowd in an unsupervised manner.
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) (Casper et al., 2023) is a promising approach for learning intelligent behaviors in the absence of a known numerical reward signal, where humans typically answer pairwise preference queries of the form “Do you prefer A or B?” (Christiano et al., 2017; Lee et al., 2021b). RLHF methods have shown success in domains from robotics (Wilde et al., 2021; Torne et al., 2023; Lee et al., 2021a) to training large language models (LLMs) (Ouyang et al., 2022). Yet, most current RLHF methods treat the preference feedback as if it came from a single human user, even if it is actually given by multiple users with potentially different backgrounds and levels of expertise. In contrast, aggregating crowdsourced data in a manner that recognizes differences between individuals could enable algorithms to analyze disagreement patterns across the crowd, detect minority groups, mitigate model bias and ensure fairness (Ouyang et al., 2022), and assist users with diverse preferences or varying degrees of noise in their feedback, whether through personalization or by learning behaviors that balance various users’ needs. Indeed, aggregating information across a crowd can result in better decisions than those of any single crowd member (Surowiecki, 2005); this has been demonstrated across many domains (Prelec et al., 2017) as the “wisdom of crowds” phenomenon.
We consider RL from human pairwise preference feedback that is crowdsourced, i.e., where (1) each data point is labeled by multiple human users and (2) these labels are aggregated to form an ensemble label (or, if applicable, identification of minority feedback groups with their ensemble labels) for downstream training. Notably, several recent RLHF works propose methods to learn from multi-user feedback, e.g. Siththaranjan et al. (2023); Poddar et al. (2024); Chakraborty et al. (2024). However, these prior approaches do not leverage crowdsourcing methods (Chen et al., 2013; Li et al., 2016) that explicitly model information from different users by incorporating annotator identifiers in the learning process; such methods could potentially augment RLHF methods with interesting properties, for example characterizing user accuracies, improving data requirements with respect to identifying multiple viewpoints, efficiently assigning users to inferred clusters based on viewpoint, and tailoring queries to specific users. In addition, while there are datasets available that study diverse preferences in multi-human feedback (Bai et al., 2022; Durmus et al., 2024; Nakano et al., 2022), these typically do not contain annotator identifiers together with data labeled by multiple annotators, for example due to cost reasons (Ouyang et al., 2022), potentially inhibiting the ability to learn and incorporate diverse preferences among human users.
@@
Figure 1: Crowd-PrefRL framework for training RL agents via crowd feedback. We assume a crowd, or ensemble, of users is queried for their preferences over pairs of segmented behaviors (A vs. B). These crowd preferences are then used to learn aggregate preference labels and reward function(s) for preference-based RL. Note that crowds might consist of users who provide diverse preference feedback according to different objectives (blue vs. orange), which we assume are unknown a priori.
This work studies preference-based RL leveraging crowd feedback in the online RL scenario, where the RL algorithm has environment access and can sample new preference pairs for the crowd to label throughout model training. Unlike prior methods, we propose to perform label aggregation via the Spectral Meta-Learner (SML) (Parisi et al., 2014) to learn noise-robust aggregate preference labels to train an aggregate reward. This approach, which leverages annotator identifiers, requires minimal modeling assumptions, is straightforward to compute, and has no hyperparameters to tune. Furthermore, it is robust to multiple viewpoints and as a byproduct, yields weights that enable principled inference of user accuracies and direct clustering of users via their inferred objectives. Specifically:
@@
    We introduce a conceptual RL framework, Crowd-PrefRL, that integrates online preference-based RL with techniques from unsupervised crowdsourcing to train AI agents with feedback from crowds of unknown expertise and reliability. Crowd-PrefRL aggregates the user feedback to learn crowdsourced preference labels, and then uses those labels for reward function learning in preference-based RL (see Figure 1). Crowd-PrefRL can also detect the presence of multiple, e.g. minority, viewpoints among the crowd, given only the crowd’s binary preference feedback.
@@
    We perform exploratory experiments in robot locomotion tasks suggesting that RL agents trained with crowd preferences outperform RL agents trained via majority-vote preferences or preferences from any individual crowd member, in particular when the spread of user errors among the crowd is large. We also show a proof-of-concept for the ability of our method to identify minority viewpoints and cluster users by viewpoint, enabling downstream personalization to diverse users.
2 Related Work
RLHF. While RL from pairwise preference queries has demonstrated numerous successes (Christiano et al., 2017; Lee et al., 2021b; a; Ouyang et al., 2022), much prior work assumes all data comes from a single human user, in some cases treating multi-user data as if it came from just one person, e.g. as in Ouyang et al. (2022). In addition, existing publicly available datasets for studying diverse human preferences omit information about annotator identifiers, e.g. Bai et al. (2022); Nakano et al. (2022); Durmus et al. (2024). Unlike these works, our approach explicitly models feedback from multiple users with different degrees of noise and who potentially give feedback with respect to multiple reward functions. Our label aggregation approach could be paired together with any RLHF method, including methods that do not learn a reward function (Rafailov et al., 2023).
Label Aggregation. Previous crowdsourcing literature in classification settings aggregates labels across users to learn crowd labels and detect patterns of outlier feedback  (Chen et al., 2013; Li et al., 2016). Parisi et al. (2014) develop the SML label aggregation method, which is robust to both noisy users and to the presence of malicious user groups. Yet, the application of label aggregation to both preference-based learning and to RL remains understudied. While Brown et al. (2019) aggregate labels via a majority vote in an imitation learning setting, unlike the SML, a majority vote cannot account for patterns across users, e.g. users who consistently provide poor feedback. In contrast, the SML performs a spectral analysis of the covariance matrix of the crowd labels to estimate crowd worker reliabilities and can identify the presence of minority user groups in the crowd.
RL from Multi-Human Feedback. Several recent works (Zhang & Kashima, 2023; Siththaranjan et al., 2023; Poddar et al., 2024) use data collected from multiple humans for preference-based reward learning. However, these works are limited to the offline RL setting, which potentially limits the expressiveness of the learned reward to behaviors in the offline dataset. Furthermore, Siththaranjan et al. (2023) and Poddar et al. (2024) infer a distribution over learned rewards across users, potentially requiring restrictive modeling assumptions and large quantities of data. E.g., Siththaranjan et al. (2023) propose two choices of reward distribution, while  Poddar et al. (2024) leverage variational inference to infer the reward distribution, incurring higher data requirements compared to the SML (7,500 preference pairs per user). In contrast, our method performs on-policy RLHF and learns aggregate preference labels to train an aggregate reward function. We utilize the SML as a label aggregation method, leveraging annotator identifiers unlike prior RLHF works. In addition, our method provides a principled way to infer user clusters and to infer each user’s accuracy with respect to either the majority reward or a particular user cluster’s reward. Furthermore, the SML is straightforward to compute, does not have any hyperparameters to tune, and is robust to the presence of multiple groups of users with differing objectives.
Several of these works pursue pluralistic preference alignment (Poddar et al., 2024; Chakraborty et al., 2024; Ramesh et al., 2024), i.e., to train AI models such as LLMs to generate responses that balance a diversity of human viewpoints. Yet, these methods often require modeling assumptions such as a specified number of user clusters or set of user beliefs (Chakraborty et al., 2024; Yao et al., 2025; Ramesh et al., 2024; Zhao et al., 2024). Also, they typically do not provide flexibility in how to balance the multiple user objectives, e.g. via allowing multiple objective weightings. For instance, Chakraborty et al. (2024); Ramesh et al. (2024) attempt to maximize the minimum utility over all individuals. Unlike these works, our approach enables user cluster inference, which could facilitate a range of downstream methods for balancing among the learned rewards.
@@ 3 Problem Formulation
In this paper, we are concerned with answering the following questions: (1) How can we effectively combine preference feedback from multiple users of unknown expertise and reliability in the absence of ground-truth information? (2) Can we learn a crowdsourced reward function that yields better performance than any individual user’s reward function?, and (3) Can we detect the presence of minority feedback (as opposed to purely noisy feedback) in the crowd and cluster users by their objectives using only the preference feedback? This section formulates the problem more precisely.
@@
We consider an RL agent that takes actions and interacts with an environment (Sutton & Barto, 2018), such that at each timestep $t$, the agent receives a state $\mathbf{s}_{t}$ from the environment and chooses an action $\mathbf{a}_{t}$ based on its policy $\pi$. In traditional episodic RL, the environment also returns a reward $r(\mathbf{s}_{t},\mathbf{a}_{t})$, and the agent seeks to maximize the total discounted sum of rewards over an episode. However, for many complex tasks, it is difficult to construct a suitable reward function. Preference-based RL approaches instead query a human user to obtain preference feedback over segments of agent behaviors and use this feedback to learn a reward function (Christiano et al., 2017; Lee et al., 2021b). A segment is defined as a length-$H$ sequence of observations and actions, $\{(\mathbf{s}_{1},\mathbf{a}_{1}),...,(\mathbf{s}_{H},\mathbf{a}_{H})\}$. Given a pair of segments $x=(A,B)$, a user indicates which segment is preferred: $+1$ in the event that $A$ is preferred over $B$ (denoted by $A\succ B$) or $-1$ if $B$ is preferred over $A$ (denoted by $A\prec B$). Preference-based RL aims to train an agent to perform human-desirable behaviors according to the human user’s preferences.
@@
Our problem formulation assumes access to a crowd, or ensemble, of users, who each assign pairwise preference labels to a sequentially-selected set $\mathcal{S}=\{x_{k}\}_{k=1}^{S}$, $S=|\mathcal{S}|$ of segment pairs. Each query $x_{k}$ takes the form $x_{k}=(A_{k},B_{k})$, where $A_{k},B_{k}$ belong to the set $\mathcal{X}$ of available segments. In addition, we assume that a majority (e.g. at least $51\%$) of the crowd provides preference feedback according to a shared understanding of desired agent behavior. More concretely, we assume the existence of an unknown ground-truth reward $r$, and that for a majority of users, their preferences are (noisily) generated with respect to $r$. Note that this formulation allows one or more minority groups of users to potentially provide feedback according to a different reward function(s). Beyond this, we do not assume any further knowledge about the crowd (e.g. expertise, reliability). Our goal is to learn a crowd-informed preference label for each segment pair using only the preference labels given by the crowd. Formally, let ${\{f_{i}\}_{i=1}^{M}}$ represent $M$ users of unknown reliability and error rate, where each user $i$ provides a preference label $f_{i}(x_{k})\in\{-1,1\}$ on each input query $x_{k}\in\mathcal{S}$. Then, let $\bm{y}$ be the vector of (unknown to the algorithm) ground-truth labels according to the true reward $r$ (i.e., in which segments with higher ground-truth reward are preferred), $\bm{y}=[y_{1},...,y_{S}]^{T}$. Using only the preference labels from the $M$ users on the segment pairs in $\mathcal{S}$ and without access to any ground-truth rewards or preferences, we seek an optimal estimate $\hat{\bm{y}}=[\hat{y}_{1},...,\hat{y}_{S}]^{T}$ of the true labels $\bm{y}$ given the crowd preference labels. We also investigate the ability to identify minority groups of users who provide feedback that consistently differs from that of the majority.
@@ 4.1 Simulating Diverse Crowds
To systematically evaluate the impact of user error rates on learning from crowdsourced user preferences, we apply the Stochastic Preference Model (Lee et al., 2021a) to simulate users with mixtures of different irrationality levels:
$P[A\succ B;\beta,\gamma]=$ $\hskip 14.22636pt\frac{e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{A},\mathbf{a}_{t}^{A})\right)}}{e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{A},\mathbf{a}_{t}^{A})\right)}+e^{\left(\beta\sum_{t=1}^{H}\gamma^{H-t}r(\mathbf{s}_{t}^{B},\mathbf{a}_{t}^{B})\right)}},$ (1)
where $\gamma\in(0,1]$ is a discount factor to model myopic behavior, and $\beta$ is a rationality constant. Note that user feedback is fully rational and deterministic as $\beta\rightarrow\infty$, whereas $\beta=0$ produces uniformly random choices. An additional parameter $\epsilon$ models users making mistakes due to accidental errors. For our analysis, each user is simulated with a random draw of parameters from the following ranges: $\gamma\in[0.98,1.0]$, $\beta\in[0.1,10.0]$, and $\epsilon\in[0,0.2]$. These parameter ranges were chosen to produce diverse user error rates primarily between $15\%$ and $40\%$, which is approximately the range of human error rates observed in a prior study of human feedback in preference-based RL ($12\%-37\%$ in Brown et al. (2019)). We then sample crowds of $M\in\{7,11,15\}$ users according to this model to study the effect of crowd size on crowd label learning.